# Multimodal Outer Arithmetic Block Dual Fusion of Whole Slide Images and Omics Data for Precision Oncology

### v1

@@ 6 Conclusion
In conclusion, the MOAD-Net framework demonstrates superior accuracy and robustness across evaluation metrics, outperforming other methods in both subtyping and survival prediction tasks. This performance highlights the unique advantages of its dual fusion strategy, combining early and late fusion to maximize predictive power and interoperability. Early fusion, by integrating omics and WSI data at the initial stages, enhances interpretability and improves classification performance through the use of ABMIL, which effectively assigns attention to crucial regions. Late fusion intermingling enriches the model’s overall representation, allowing MOAD-Net to capture more complex relationships between modalities, ultimately enhancing performance across all metrics. The outer arithmetic operations—addition, subtraction, multiplication, and division—further strengthen this dual fusion approach, enabling MOAD-Net to produce a richer, more informative representation. As a result, ABMIL-MOAD-Net excels in both accuracy and consistency, making it well-suited for complex multimodal predictive tasks in precision oncology.

### v2

@@ 5 Conclusion
The MOAD-FNet framework advances CNS tumor subtyping by effectively integrating DNA methylation and WSI data through a novel dual fusion approach. Our framework addresses the under-explored potential of combining these modalities, with the potential for distinguishing subtle morphological differences between tumor subtypes from DNA methylation profiling. The early fusion stage, implemented through an MLP-based mapping of WSI and methylation features, enables interpretable visualization while maintaining low dimensionality. The late fusion stage, enriched by outer arithmetic operations with MOAB, captures complex inter-modal relationships beyond simple addition. MOAD-FNet demonstrates superior performance across all evaluation metrics and exhibits robust scalability with different architectures. The framework’s consistent success in both tumor subtyping and survival prediction establishes its practical utility for precision oncology. These results highlight how integration of multiple data modalities can enhance diagnostic accuracy while preserving clinical interpretability, thereby advancing automated approaches to CNS tumor classification.